a particle is projected from point A with speed u and angle of projection is 60 degree at some instant magnitude of velocity of particle is v and it makes an angle theata (o) with horizontal . if radius of curvature of path of particle at the given instant is 8/3(3)^1/2times the minimum radius of curvature during the whole flight then choose incorrect options ,
a)theata=37 degree
b)theata=30 degree
c)v=u/2
d)v=u/2(3)^1/2

Grade:12

1 Answers

Akash Kumar Dutta
11 years ago

Dear Prakhar,
Let (o) represent theta.
Radius of curvature is given by=>R=v^2/gcos (o).
and vcos(o)=u/2.
R(min)=u^2/4g
using these three eq. we get (o)=30 degree.
Hence (a) option is wrong
And v=u/3^1/2
Hence option(c),(d) are also wrong.
The ans is...(a),(c),(d)
ANS

Regards.